How many coulontos of charge move across a 120 V source when 1440 J of energy is

required?


1
Expert's answer
2022-01-29T02:48:53-0500

The energy, voltage and charge are related as


"\Delta E = q\cdot \Delta V 
\\ \text{}
\\ \therefore q = \frac{\Delta E }{ \Delta V}= \frac{1440\,J }{ 120\,V}= 12 \frac{\cancel{J }}{ \cancel{J}/C}= 12\,C"


Therefore, the amount of charge moved is 12 Coulombs.